OpenStudy (anonymous):

I need help solving a problem 15x^4-11x^2+2=0

OpenStudy (anonymous):

\[x = ±\sqrt{\frac{2}{5}} \approx ±0.632456\] \[x = ±\frac{1}{\sqrt{3}}\approx±0.57735\]

OpenStudy (anonymous):

Let \[z=x^2\] The equation is now \[15z^2 - 11z +2=0\] Use the Quadratic Formula to find the roots of the equation: \[z = (-b \pm \sqrt{b^2-4ac}) / 2a\] Then, the solution is \[x=\pm \sqrt{z}\]

OpenStudy (anonymous):

Using the above method, the roots are: \[\pm \sqrt{10}/5, \pm \sqrt{3}/3\]